Which is greater in each of the following:  2/3, 5/2

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

The LCM of the denominators 3 and 2 is 6
(2/3) = [(2 × 2)/ (3 × 2)] = (4/6)
And (5/2) = [(5 × 3)/ (2 × 3)] = (15/6)
Now,
4 < 15
So, (4/6) < (15/6)
∴2/3 < 5/2
Hence, 5/2 is greater.

​​Thanks